# Efficient Exact Algorithms for Minimum Covering of Orthogonal Polygons with Squares

### v1

@@ 5.2 Recursion with Separating Squares
Therefore in all cases, we can map knobs of $Q\cup S$ for any $Q\in\mathcal{Q}$ to distinct knobs in $P$. Therefore the total number of knobs of $Q\cup S$ can be at most $k$, if $P$ had at most $k$ knobs.
Using the following framework, we find a non-knob convex vertex $v$ (if any) in $\mathcal{O}(n)$ time (Remark 5.4). We construct $S:=MCS(v)$, construct $\mathcal{Q}$ and use this recursion to recurse into $|\mathcal{Q}|$ similar instances where the number of knobs do not increase. We observe that each recursive step can also be done in linear time by a simple traversal of the polygon vertices in order.
@@
If $P$ is an orthogonal polygon with $n$ vertices and at most $k$ knobs, we can either report that no non-knob convex vertex exists, or $z\leq 12$ new smaller instances of orthogonal polygons $P_{1},\ldots P_{z}$ which individually have at most $k$ knobs, in $\mathcal{O}(n)$ time.
@@
Proof 5.13.
Let $n_{x},n_{v}$ be the number of convex vertices and concave vertices in $P$ respectively. We must $n_{v}=n_{x}-4$ because $P$ is an orthogonal polygon with no holes. Moreover, as all convex vertices are part of knobs and there are at most $k$ knobs (containing two distinct vertices), we must have $n_{x}\leq 2k$. Therefore the total number of vertices is $n_{x}+n_{v}=2n_{x}-4\leq 4k-4$.
Therefore, for such polygons, we have $n=\mathcal{O}(k)$. We can detect if this is the case (i.e. $P$ does not have a non-knob convex vertex) in $\mathcal{O}(n)=\mathcal{O}(k)$ time (Lemma 5.11) and apply Algorithm 4 to solve OPCS $(P)$ in time $\mathcal{O}(n^{14})=\mathcal{O}(k^{14})$ (Theorem 4.17).
@@
Given an orthogonal polygon $P$ with $n$ vertices and at most $k$ knobs and with no non-knob convex vertices, we can solve OPCS in $\mathcal{O}(k^{14})$ time.
@@ 5.3 A Recursive Algorithm
With the previous results in hand, we can construct an exact algorithm, that solves OPCS on an arbitrary orthogonal polygon $P$ with $n$ vertices and at most $k$ knobs.
The algorithm follows the framework of (i) recurse as in Lemma 5.11 whenever possible and (ii) detect and solve base cases as in Lemma 5.14. We formally state the algorithm as follows.
@@
if $P$ has no non-knob convex vertex then $\triangleright$ $\mathcal{O}(n)$ from Lemma 5.11
     $C\leftarrow$OPCS $(P)$triangleright using Algorithm 4, $\mathcal{O}(k^{14})$ time from Lemma 5.14
@@
$v\leftarrow$ some non-knob convex vertex $\triangleright$ $\mathcal{O}(n)$ from Lemma 5.11
$S\leftarrow MCS(v)$triangleright $\mathcal{O}(n)$ from Lemma 3.2
Construct $\mathcal{Q}$ as in Lemma 5.6 $\triangleright$ total $\mathcal{O}(n)$ time
@@
for $Q\in\mathcal{Q}$ do $\triangleright$ recurse, Algorithm 5, $|\mathcal{Q}|\leq 12$
@@
return $(s-(|\mathcal{Q}|-1))$triangleright application of Lemma 5.6
5.4 Analysis of Separating-Square-Recursion
@@
Let an orthogonal polygon $P$ with $n$ vertices and at most $k$ knobs be the original input to the topmost level invocation of Algorithm 5. At some recursive step, let the input be $P_{0}$, and let the algorithm choose $v$ to be a non-knob convex vertex of $P_{0}$. Then $v$ must also be a non-knob convex vertex of the original polygon $P$. Moreover, any corner of $S$ that is also a vertex of $Q\cup S$ can not be a non-knob convex vertex.
Proof 5.16.
@@
Next, if $v$ is a vertex participating in a knob $(u,v)$ in $P$, then $u$ is a convex vertex. Due to this, at any intermediate recursion step, there cannot a polygon $P^{\prime}$ ($P_{0}$ in particular) with $v$ as a vertex and a concave endpoint to the edge originating from $v$ and along $uv$. Thus, $v$ must be a non-knob convex vertex in $P$.
We now prove that if $v$ is chosen as a non-knob convex vertex at some recursive step, then no subsequent recursive steps can again choose $v$ to construct their separating square.
@@
Proof 5.20.
Due to Lemma 5.17, the number of recursive steps where Algorithm 5) recurses (i.e. internal nodes in recursion tree), is bounded by the number $n$ of vertices in the original input polygon. As any recursive step calls at most $12$ more recursive steps, the number of steps where Algorithm 5 achieves the base-case condition (no non-knob convex vertex found) and calls Algorithm 4 is bounded by $12n$ (i.e. leaf nodes in recursion tree). Therefore we have the above result.
Finally, we complete our algorithm with bounding the running time.
@@
Separating-Square-Recursion (Algorithm 5) when run on an arbitrary orthogonal polygon $P$ with $n$ vertices and at most $k$ knobs, outputs the value of OPCS $(P)$ in $\mathcal{O}(n^{2}+k^{14}\cdot n)$ time.
Proof 5.22.
Each internal node of the recursion tree for the algorithm takes $\mathcal{O}(n)$ time. Therefore, following from the bound in Lemma 5.19, the total time taken by the internal nodes of the recursion tree for the algorithm is $\mathcal{O}(n^{2})$.
Each leaf node of the recursion tree is an execution of Algorithm 4, each taking $\mathcal{O}(k^{14})$ time (Lemma 5.14). Now, as there are $\mathcal{O}(n)$ leaf nodes in the recursion tree (Lemma 5.19), the base conditions together take $\mathcal{O}(k^{14}\cdot n)$ time.
Hence, Separating-Square-Recursion runs in $\mathcal{O}(n^{2}+k^{14}\cdot n)$ time.
@@
Note that we should only prefer Algorithm 5 when $k=o(n^{13/14})$, otherwise Algorithm 4 provides better or same asymptotic run time.
@@
If there is an algorithm solving OPCS in time $T(n)$ for polygons with $n$ vertices, there is also another algorithm solving $p$-OPCS on polygons with $n$ vertices and at most $k$ knobs in time $\mathcal{O}(n^{2})+n\cdot T(4k-4)$.
@@ 5.5 Discussion on Orthogonally Convex Polygons
  • $P$ is a simply connected polygon with polygon edges parallel to either the x-axis or the y-axis.
  • The intercept of any line parallel to x-axis or y-axis with $P$ produces one continuous (possibly empty) line segment parallel to x-axis or y-axis.
@@
It is easy to construct a simple orthogonal polygon having arbitrarily large number of knobs. However, we show that for orthogonally convex polygons, the number of knobs must be exactly $4$.
@@
Proof 5.27.
Existence. Consider the leftmost vertical line that intersects $P$ to form a non-empty vertical line segment of intersection. This must intersect with a vertical polygon edge of $P$ (otherwise the vertical line can be moved more to the left). Clearly these endpoints of this vertical polygon edge form a left knob (otherwise the vertical line can again be moved more to the left). Hence a left knob always exists. Symmetric arguments yield that a top knob, a bottom knob and a right knob exists as well.
@@
Uniqueness. We will show that there cannot be two left knobs implying that there is exactly one left knob. For the sake of contradiction, consider that there are two left knobs $(v_{i},v_{i+1})$ and $(v_{j},v_{j+1})$ with $x_{i}=x_{i+1}\geq x_{j}=x_{j+1}$ (refer to Figure 14). Consider the vertical line $x=x_{i}$. This line intersects $P$ at the entire edge $(v_{i},v_{i+1})$ with the intercept the y coordinates being $[\min(y_{i},y_{i+1}),\max(y_{i},y_{i+1})]$. Now consider any curve lying inside $P$ with $v_{i}$ and $v_{j}$ as endpoints. This curve must also intersect the line $x=x_{i}$ (as $x_{j}\leq x_{i}$) at some y coordinate outside $[\min(y_{i},y_{i+1}),\max(y_{i},y_{i+1})]$. This means the intersection of the line $x=x_{i}$ with $P$ is not a single line segment, which contradicts the assumption that $P$ is orthogonally convex. Therefore, there can be exactly one left knob. Subsequently, by symmetric arguments this would mean that there is exactly one right knob, exactly one top knob and exactly one bottom knob.
Therefore, if we use Algorithm 5 to solve OPCS on orthogonally convex polygons, we can substitute $k=4$ in the analysis.
@@
As orthogonally convex polygons have exactly $k=4$ knobs, Separating-Square-Recursion (Algorithm 5) takes $\mathcal{O}(n^{2})$ time to solve OPCS on orthogonally convex polygons.
@@ 6 Hardness Results for Polygons with Holes
In this Section, we consider the OPCSH problem. We consider the setting as described in [3], i.e. we need to solve OPCSH in orthogonal polygons, where the input consists of all $N$ lattice points lying in the polygon or on the boundary. In [3], a proof is given to show that OPCSH is NP-complete if the input size is polynomial in $N$. In this section, we fix an issue with the proof and state other results obtained as a consequence.
@@ 6.1 Issue with the existing proof
The proof in [3] for NP-completeness of OPCSH gives a reduction from Planar 3-CNF [8]. The reduction gives a polynomial time algorithm to reduce any Planar 3-CNF instance (say $\psi$) and transforms it to an instance of OPCSH. First the formula is negated to obtain $\phi=\neg\psi$, which is a formula in disjunctive normal form (DNF) due to De-Morgan’s law. This means that $\psi$ is satisfied if and only if $\phi$ evaluates to false on some truth assignment of the variables. Now, $\phi$ is reduced to an instance of OPCSH using three kinds of gadgets: wires, variable gadgets and junction gadgets. A variable gadget is introduced for each variable of $\phi$ and a junction gadget is introduced for each conjunction in $\phi$. Further wire gadgets (or simply wires) are introduced to ‘connect’ junction gadgets to variable gadgets.
@@
The variable gadgets, wires and the junction gadgets shall be placed according to the formula $\phi$. For each variable gadget or each wire, it is easy to compute in polynomial time the minimum number of squares needed to cover the gadget. It is possible that some of the squares used for covering a wire may cover a region inside a junction gadget. Each variable gadget permits two kinds of covering (representing the two kinds of truth assignment to a variable in $\phi$). Let us denote these coverings by true-covering and false-covering, respectively - both coverings require the same number of squares. Similarly we can extend this observation to see that there are two ways in which a variable gadget and its adjacent wires can be covered optimally - and both coverings use the same number of squares. A combination of how a variable gadget is covered affects which portions of all its corresponding junction gadgets (i.e. gadgets for the and-clauses containing the corresponding variable in $\phi$) will be covered. We will keep the construction of variable gadgets and wires unchanged from the construction in [3]. Let $V$ be the number of squares required to cover all variable gadgets and $W$ be the number of squares required to cover all wire gadgets.
@@
In [3], the authors claim that the junction gadgets (or simply, a junction) are such that, the minimum number of squares needed to cover a junction gadget is $12$ if all three variable gadgets corresponding to the and-clause are covered with a covering which makes the and-clause evaluate to true (i.e. true-covering if the variable appears non-negated in the clause, or false-covering if the variable is negated in the clause); otherwise the minimum number of squares needed is $13$. Therefore, the junctions act similar to and-gates. Now, if the total number of junctions be $j$, then [3] concludes by stating that $\psi$ is satisfiable if and only if the reduced instance can be covered with less than $(V+W+13j)$ squares; otherwise exactly $(V+W+13j)$ squares are required. Hence solving OPCSH would also solve Planar 3-CNF.
@@
However, we take a deeper look into this reduction by performing a case work on the evaluated value of $\psi$ for possible truth assignments of variables.
@@
  • Case II, $\psi$ is false for all assignments. This implies $\phi=\neg\psi$ is true for all assignments. Since $\phi$ is in DNF, all assignments must render at least one and-clauses as true. Therefore, for any arbitrary minimum covering of the variable gadgets, there would be at least one junction that can be covered with $12$ squares (while others take at most $13$ squares). This means the minimum number of squares needed to cover such an instance is less than $V+W+13j$.
  • Case III, $\psi$ is true for some assignments and false for some. This implies $\phi=\neg\psi$ is also true for some assignments and false for some. Consider any such assignment such that $\phi$ is true and cover the variable clauses accordingly. Since $\phi$ is in DNF, this assignments must render at least one and-clauses as true. Therefore, if we cover the variable gadgets corresponding to such an assignment, there would be at least one junction that can be covered with $12$ squares (while others take at most $13$ squares). This means the minimum number of squares needed to cover such an instance is less than $V+W+13j$.
Therefore, the minimum number of squares to cover the reduced instance is equal to $V+W+13j$ if and only if $\psi$ is true for all assignments (Case I) and less than $V+W+13j$ otherwise. Therefore, solving OPCSH solves tautology for $\psi$ instead of satisfiablity. Moreover, tautology in a CNF formula can be checked in linear time (Since $\psi$ is a tautology if and only if for all clauses $c$ in $\phi$ there is a variable $x$ such that both $x$ and $\neg x$ appear in $c$). Hence this reduction does not prove NP-hardness of OPCSH.

### v2

@@ 6.2 Recursion with separating squares
Therefore in all cases, we can map knobs of $Q\cup S$ for any $Q\in\mathscr{Q}$ to distinct knobs in $P$. Therefore the total number of knobs of $Q\cup S$ can be at most $k$, if $P$ had at most $k$ knobs.
We use the following framework:
  1. 1.
    Find a non-knob convex vertex $v$ (if any) in $\mathcal{O}(n)$ time (Remark 6.4).
  2. 2.
    If no such non-knob convex vertex exists, solve OPCS using Algorithm 3 (base cases).
  3. 3.
    If it exists, construct $S:=MCS(v)$, construct $\mathscr{Q}$ and use this recursion to recurse into $|\mathscr{Q}|$ similar instances where the number of knobs do not increase.
Each recursive step which is not a base case, can be done in linear time $\mathcal{O}(n)$, by a simple traversal of the polygon vertices. As $|\mathscr{Q}|\leq|\mathscr{Q}^{\prime}|\leq 12$, we get the following result.
@@
If $P$ is an orthogonal polygon with $n$ vertices and at most $k$ knobs, in $\mathcal{O}(n)$ time, we can either report that no non-knob convex vertex exists, or $z\leq 12$ smaller instances of orthogonal polygons $P_{1},\ldots,P_{z}$ which individually have at most $k$ knobs and $n$ vertices.
@@
Proof 6.12.
Let $n_{x},n_{v}$ be the number of convex vertices and concave vertices in $P$ respectively. Since $P$ is an orthogonal polygon with no holes, we have $n_{v}=n_{x}-4$. Moreover, as only convex vertices are part of knobs and there are at most $k$ knobs, we must have $n_{x}\leq 2k$. Therefore the total number of vertices is $n_{x}+n_{v}=2n_{x}-4\leq 4k-4$.
Therefore, for such polygons, we have $n=\mathcal{O}(k)$. We can detect this in $\mathcal{O}(n)=\mathcal{O}(k)$ time, and solve OPCS in $\mathcal{O}(n^{10})=\mathcal{O}(k^{10})$ time using Algorithm 3.
@@
Given an orthogonal polygon $P$ with $n$ vertices and at most $k$ knobs and with no non-knob convex vertices, we can solve OPCS in $\mathcal{O}(k^{10})$ time.
@@ 6.3 A recursive algorithm
We now have the results to design our exact algorithm solving $p$-OPCS on input orthogonal polygons $P$ having $n$ vertices and at most $k$ knobs. We formally state our recursive framework as an algorithm.
@@
if $P$ has no non-knob convex vertex then $\triangleright$ $\mathcal{O}(n)$, Lemma 6.10
     $C\leftarrow$OPCS $(P)$triangleright Algorithm 3, $\mathcal{O}(k^{10})$, refer to Lemma 6.13
@@
$v\leftarrow$ some non-knob convex vertex
$S\leftarrow MCS(v)$triangleright $\mathcal{O}(n)$, Lemma 4.1
Construct $\mathscr{Q}$ as in Lemma 6.6 $\triangleright$ $\mathcal{O}(n)$
@@
for $Q\in\mathscr{Q}$ do $\triangleright$ recurse, Algorithm 4, $|\mathscr{Q}|\leq 12$
@@
return $(s-(|\mathscr{Q}|-1))$triangleright Lemma 6.6
@@
Let an orthogonal polygon $P$ with $n$ vertices and at most $k$ knobs be the original input to Algorithm 4. At some recursive step, let the input be $P_{0}$, and let the algorithm choose $v$ to be a non-knob convex vertex of $P_{0}$. Then $v$ must also be a non-knob convex vertex of the original polygon $P$. Moreover, any corner of $S$ that is also a vertex of $Q\cup S$ can not be a non-knob convex vertex.
Proof 6.15.
@@
Next, if $v$ is a vertex participating in a knob $(u,v)$ in $P$, then $u$ is a convex vertex. Due to this, at any intermediate recursion step, there cannot a polygon $P^{\prime}$ ($P_{0}$ in particular) with $v$ as its vertex, that has a concave endpoint to the edge originating from $v$ and along $uv$. Thus, $v$ must be a non-knob convex vertex in $P$.
We now prove that if $v$ is chosen as a non-knob convex vertex at some recursive step, then no subsequent recursive steps can again choose $v$.
@@
Proof 6.19.
Due to Lemma 6.16, the number of recursive steps where Algorithm 4 recurses (internal nodes in recursion tree), is bounded by the number $n$ of vertices in the original input polygon. As any recursive step calls at most $12$ more recursive steps, the number of steps where Algorithm 4 achieves the base-case condition and calls Algorithm 3 is bounded by $12n$ (leaf nodes in recursion tree).
Finally, we complete the analysis of our algorithm by bounding the total running time.
@@
Separating-Square-Recursion (Algorithm 4) when run on an orthogonal polygon $P$ with $n$ vertices and at most $k$ knobs, solves OPCS in $\mathcal{O}(n^{2}+k^{10}\cdot n)$ time.
Proof 6.21.
Each internal node of the recursion tree for the algorithm takes $\mathcal{O}(n)$ time (Lemma 6.10). Therefore, following from the bound in Lemma 6.18, the total time taken by the internal nodes of the recursion tree for the algorithm is $\mathcal{O}(n^{2})$.
Each leaf node of the recursion tree is an execution of Algorithm 3, each taking $\mathcal{O}(k^{10})$ time (Lemma 6.13). Now, as there are $\mathcal{O}(n)$ leaf nodes in the recursion tree (Lemma 6.18), the base cases together take $\mathcal{O}(k^{10}\cdot n)$ time. Total running time becomes $\mathcal{O}(n^{2}+k^{10}\cdot n)$.
@@
We should only prefer Algorithm 4 when $k=o(n^{9/10})$, otherwise Algorithm 3 provides better or the same asymptotic running time.
@@
If there is an algorithm solving OPCS in time $T(n)$ for polygons with $n$ vertices, there exists an algorithm solving $p$-OPCS on orthogonal polygons with $n$ vertices and at most $k$ knobs in time $\mathcal{O}(n^{2})+n\cdot T(4k-4)$.
@@ 6.4 Discussion on orthogonally convex polygons
  • $P$ is a simply connected polygon with polygon edges parallel to the x-axis or the y-axis.
  • The intercept of any line parallel to x-axis or y-axis with $P$ produces one continuous (possibly empty) line segment.
@@
It is easy to construct a simple orthogonal polygon having an arbitrarily large number of knobs. However, we show that for orthogonally convex polygons, the number of knobs must be exactly $4$.
@@
Proof 6.26.
Existence. Consider the leftmost vertical line that intersects $P$ to form a non-empty vertical line segment of intersection. This must intersect with a vertical polygon edge of $P$. The endpoints of this vertical polygon edge form a left knob. Hence a left knob always exists. Symmetric arguments yield that a top knob, a bottom knob and a right knob exists as well.
@@
Uniqueness. We show that there cannot be two distinct left knobs. For the sake of contradiction, consider that there are two left knobs $(v_{i},v_{i+1})$ and $(v_{j},v_{j+1})$ with $x_{i}=x_{i+1}\geq x_{j}=x_{j+1}$ (refer to Figure 16). Consider the vertical line $x=x_{i}$. This line intersects $P$ at the entire edge $(v_{i},v_{i+1})$ with the intercept the y coordinates being $[\min(y_{i},y_{i+1}),\max(y_{i},y_{i+1})]$. Now consider any curve lying inside $P$ with $v_{i}$ and $v_{j}$ as endpoints. This curve must also intersect the line $x=x_{i}$ (as $x_{j}\leq x_{i}$) at some y coordinate outside $[\min(y_{i},y_{i+1}),\max(y_{i},y_{i+1})]$. This means the intersection of the line $x=x_{i}$ with $P$ is not a single line segment, which contradicts the assumption that $P$ is orthogonally convex. Therefore, there can be exactly one left knob. By symmetric arguments, there is exactly one right knob, exactly one top knob and exactly one bottom knob.
Therefore, if we use Algorithm 4 to solve OPCS on orthogonally convex polygons, we can substitute $k=4$ in the analysis, giving us a running time of $\mathcal{O}(n^{2}+n\cdot 4^{10})=\mathcal{O}(n^{2})$.
@@
Separating-Square-Recursion (Algorithm 4) takes $\mathcal{O}(n^{2})$ time to solve OPCS on orthogonally convex polygons.
@@ 7 Hardness Results for Polygons with Holes
In this Section, we consider the OPCSH problem. We consider the setting as described by Aupperle et. al [4]: we need to solve OPCSH in orthogonal polygons, where the input consists of all $N$ lattice points lying in the polygon or on the boundary. We first state the issue with their claimed proof for OPCSH being NP-complete. Then we state a correct proof of hardness. We will use some structures defined in their paper, but use novel constructions for some gadgets used in the reduction.
@@ 7.1 Issue with the existing proof
The proof by Aupperle et. al [4] for NP-completeness of OPCSH gives a reduction from Planar 3-CNF [15] to OPCSH. The reduction gives a polynomial-time algorithm to reduce any Planar 3-CNF instance (say $\psi$) and transforms it to an instance of OPCSH. First, the formula is negated to obtain $\phi=\neg\psi$, which is a formula in disjunctive normal form (DNF) due to De-Morgan’s law. This means that $\psi$ is satisfied if and only if $\phi$ evaluates to false on some truth assignment of the variables. Now, $\phi$ is reduced to an instance of OPCSH using three kinds of gadgets: wires, variable gadgets and junction gadgets. A variable gadget is introduced for each variable of $\phi$ and a junction gadget is introduced for each conjunction in $\phi$. Further wire gadgets (or simply, wires) are introduced to ‘connect’ junction gadgets to variable gadgets.
@@
The variable gadgets, wires and the junction gadgets shall be placed according to the formula $\phi$. For each variable gadget or each wire, it is easy to compute, in polynomial time, the minimum number of squares needed to cover the gadget. It is possible that some of the squares used for covering a wire may cover a region inside a junction gadget. Each variable gadget permits two kinds of covering (representing the two kinds of truth assignment to a variable in $\phi$). Let us denote these coverings by true-covering and false-covering, respectively, Both coverings require the same number of squares. Similarly we can extend this observation to see that there are two ways in which a variable gadget and its adjacent wires can be covered optimally; both coverings use the same number of squares. A combination of how a variable gadget is covered affects which portions of all its corresponding junction gadgets (gadgets for the and-clauses containing the corresponding variable in $\phi$) will be covered. Let $V$ be the number of squares required to cover all variable gadgets and $W$ be the number of squares required to cover all wire gadgets.
@@
In the paper by Aupperle et. al [4], it is claimed that the junction gadgets (or simply, a junction) are such that, the minimum number of squares needed to cover a junction gadget is $12$ if all three variable gadgets corresponding to the and-clause are covered with a covering which makes the and-clause evaluate to true (true-covering if the variable appears non-negated in the clause, or false-covering if the variable is negated in the clause); otherwise the minimum number of squares needed is $13$. Therefore, junctions act similar to and-gates. Let the total number of junctions be $j$. Then the paper concludes by stating that $\psi$ is satisfiable if and only if the reduced instance can be covered with less than $(V+W+13j)$ squares; otherwise exactly $(V+W+13j)$ squares are required. Hence solving OPCSH would also solve Planar 3-CNF.
@@
We take a deeper look into this reduction by performing the following case work on the behaviour of $\psi$ on various truth assignments of the variables.
@@
  • Case II, $\psi$ is false for all assignments. This implies $\phi=\neg\psi$ is true for all assignments. Since $\phi$ is in DNF, all assignments must render at least one and-clauses as true. Therefore, for any arbitrary minimum covering of the variable gadgets, there would be at least one junction that can be covered with $12$ squares (while others take at most $13$ squares). This means the minimum number of squares needed to cover such an instance is strictly less than $V+W+13j$.
  • Case III, $\psi$ is true for some assignments and false for some. This implies $\phi=\neg\psi$ is also true for some assignments and false for some. Consider any assignment such that $\phi$ is true. Since $\phi$ is in DNF, this assignments must render at least one and-clauses as true. Therefore, if we cover the variable gadgets corresponding to such an assignment, there would be at least one junction that can be covered with $12$ squares (while others take at most $13$ squares). This means the minimum number of squares needed to cover such an instance is strictly less than $V+W+13j$.
Therefore, the minimum number of squares to cover the reduced instance is equal to $V+W+13j$ if and only if $\psi$ is true for all assignments (Case I) and less than $V+W+13j$ otherwise (Case II & Case III). Therefore, solving OPCSH solves tautology for $\psi$ instead of satisfiablity. Moreover, tautology in a CNF formula can be checked in linear time (Since $\psi$ is a tautology if and only if for all clauses $c$ in $\phi$ there is a variable $x$ such that both $x$ and $\neg x$ appear in $c$). Hence this reduction does not prove NP-hardness of OPCSH.